Find the general solution of these differential equations.

1) dy/dx = x/y2 2)dy/dx = x+2 / y


1
Expert's answer
2021-09-13T00:26:05-0400

"1. \frac{dy}{dx} = \frac{x}{y^2}\\
\text{Cross-multiplying we have,}\\
y^2dy= (x+2)dx\\
\text{Integrating both sides we have}\\
\frac{y^3}{3}=\frac{x^2}{2}+c\\

2. \frac{dy}{dx}=\frac{x+2}{y}\\
=ydy = (x+2)dx\\
=\frac{y^2}{2}=\frac{x^2}{2}+2x+c"